# SWA-SOP: Spatially-aware Window Attention for Semantic Occupancy Prediction in Autonomous Driving

### v1

@@ I Introduction
Recent SOP methods have adopted Transformer-based architectures to improve scene completion by leveraging attention mechanisms. However, despite their ability to capture long-range dependencies, existing Transformer-based SOP models face two critical challenges. First, they lack explicit modeling of local spatial context, which is essential for maintaining geometric consistency and accurate reasoning in 3D environments. Second, the reliance on sparse depth-guided query sampling, commonly used to reduce the high computational cost of full 3D self-attention, limits their effectiveness in distant or occluded regions where LiDAR data is extremely sparse and unevenly distributed.
@@ II-A LiDAR-based and Camera-based 3D Perception
Processing LiDAR point clouds and images into meaningful 3D representations is a core challenge in perception. For LiDAR-based semantic segmentation, a common strategy is to voxelize the point cloud and apply sparse convolutions for efficient processing [1, 2]. 3D object detection methods typically operate on either raw point clouds or voxelized representations. Point-based approaches [3, 4, 5] directly process irregular point sets to preserve geometric details, while voxel-based methods [6] convert point clouds into regular grids to enable efficient processing with 3D CNNs. Hybrid methods like CenterPoint [7] leverage sparse convolutions to balance accuracy and efficiency. In parallel, monocular and stereo depth estimation methods [8, 9, 10] aim to predict depth maps or point clouds from RGB inputs. To further enhance perception, sensor fusion techniques combine LiDAR and camera data by integrating accurate depth with rich visual semantics. For example, DeepFusion [11] aligns features from both modalities through attention-based fusion, while SLCF-Net [12] introduces a sequential fusion framework for semantic scene completion.
@@ II-B Semantic Occupancy Prediction (SOP)
Semantic Occupancy Prediction (SOP) was first formulated by Song et al. [13] as the task of inferring a complete, semantically labeled 3D occupancy grid from a single RGB-D image. With the introduction of SemanticKITTI [14], SOP has expanded to large-scale outdoor LiDAR data, motivating the development of more scalable architectures. Early approaches were primarily convolutional, such as LMSCNet [15], which applied hierarchically convolutional U-Net to 3D grids for efficiency, and JS3C-Net [2], which jointly performed segmentation and completion. Recently, SOP from RGB inputs has attracted increasing attention. MonoScene [16] pioneered the use of monocular images for SOP by projecting depth estimates into voxel space and performing volumetric completion. Building on this direction, VoxFormer [17] and OccFormer [18] further introduce transformer-based architectures, refining voxel representations with 3D deformable attention [19]. However, these methods rely on query-based sampling, which can struggle in distant or sparsely observed regions due to depth estimation errors. Besides, generative approaches such as DiffSSC [20] leverage denoising diffusion guided by segmentation features, though at the cost of increased computational complexity.
@@ III Method
We formulate semantic occupancy prediction as the task of estimating a dense semantic volume from a sparse point cloud. Given a LiDAR point cloud $\mathcal{X}=\{p_{1},...,p_{N}\}\subset\mathbb{R}^{3}$, where each $p_{i}$ denotes a 3D coordinate, the goal is to predict a semantic occupancy volume $\hat{\mathcal{Y}}$ defined over a discretized 3D space $\mathcal{V}\subset\mathbb{R}^{3}$. The space $\mathcal{V}$ is divided into a regular voxel grid of size $H\times W\times D$, and each voxel in $\hat{\mathcal{Y}}$ is assigned a label from the set =$\mathcal{L}=\{0,1,...,C\}$, where 00 indicates empty space and $1$ to $C$ correspond to semantic classes. The model aims to infer $\hat{\mathcal{Y}}\in\mathcal{L}^{H\times W\times D}$ based on the sparse input $\mathcal{X}$, including predictions for unobserved regions, to reconstruct a complete semantic representation of the scene.
@@ III-C Center Query
$q_{c}=\begin{cases}\mathrm{FF}_{q}(f_{c}),&\text{if the}f_{c}\text{is non-empty}\\ \mathrm{FF}_{q}(p_{c}),&\text{otherwise}\end{cases}$ (3)
@@ V Conclusions
In this work, we proposed SWA-SOP, a novel semantic occupancy prediction framework. By adopting a sliding-window mechanism that selectively computes attention over valid regions, SWA-SOP significantly improves the efficiency of global attention-based models. Within each window, we leverage spatial embeddings and spatial queries to perform localized attention, offering greater flexibility than traditional convolutional operations. The experiments on the SemanticKITTI benchmark demonstrate that our LiDAR-based method achieves state-of-the-art performance compared to existing approaches. Furthermore, we show that the proposed SWA module can be effectively integrated into image-based pipelines, leading to consistent performance gains. These results highlight the versatility of SWA and its potential for deployment across diverse sensor configurations in autonomous driving perception systems.

### v2

@@ I Introduction
Recent SOP methods have adopted Transformer-based architectures [1, 2, 3] to improve scene completion by leveraging attention mechanisms. However, despite their ability to capture long-range dependencies, existing Transformer-based SOP models face two critical challenges. First, they lack explicit modeling of local spatial context, which is essential for maintaining geometric consistency and accurate reasoning in 3D environments. Second, the reliance on sparse depth-guided query sampling, commonly used to reduce the high computational cost of full 3D self-attention, limits their effectiveness in distant or occluded regions where LiDAR data is extremely sparse and unevenly distributed.
@@ II-A LiDAR-based and Camera-based 3D Perception
Processing LiDAR point clouds and images into meaningful 3D representations is a core challenge in perception. For LiDAR-based semantic segmentation and object detection, common approaches include voxel-based and point-based methods. Voxel-based models [4, 5] convert point clouds into regular grids for efficient processing with 3D CNNs or sparse convolutions. Point-based methods [6, 7, 8, 9] operate directly on irregular point sets to preserve fine-grained geometry. Hybrid methods such as CenterPoint [10] combine both paradigms to balance accuracy and efficiency. In parallel, monocular and stereo methods [11, 12, 13] estimate depth or generate pseudo-LiDAR from RGB images. Sensor fusion techniques further enhance perception by combining LiDAR accuracy with visual semantics. SLCF-Net [14], for example, introduces a sequential fusion framework for semantic scene completion. Recent advances in domain-adaptive modeling [15] and prompt-driven reasoning [16] further suggest promising directions for improving robustness and transferability in multimodal perception.
@@ II-B Semantic Occupancy Prediction (SOP)
Semantic Occupancy Prediction (SOP) was first formulated by Song et al.[17] as the task of inferring a complete, semantically labeled 3D occupancy grid from a single RGB-D image. With the introduction of SemanticKITTI[18], SOP has expanded to large-scale outdoor LiDAR data, motivating the development of more scalable architectures. Early methods such as LMSCNet [19] adopted hierarchical 3D convolutions for efficient voxel processing. Recently, SOP from monocular images has attracted increasing attention. onoScene [20] initiated this direction via depth-to-voxel projection, while OC-SOP [21] builds on it by introducing object-centric awareness for enhanced semantic and geometric reasoning. In parallel, transformer-based methods [22, 23] refine voxel features using deformable attention [24], but often suffer from sparsity and errors in depth estimation, especially in distant or occluded regions. Alternatively, DiffSSC [25] explores a generative approach based on diffusion processes, offering a different modeling paradigm at the cost of higher computational complexity.
@@ III Method
We formulate semantic occupancy prediction as the task of estimating a dense semantic volume from a sparse point cloud. Given a LiDAR point cloud $\mathcal{X}=\{p_{1},...,p_{N}\}\subset\mathbb{R}^{3}$, where each $p_{i}$ denotes a 3D coordinate, the goal is to predict a semantic occupancy volume $\hat{\mathcal{Y}}$ defined over a discretized 3D space $\mathcal{V}\subset\mathbb{R}^{3}$. The space $\mathcal{V}$ is divided into a regular voxel grid of size $H\times W\times D$, and each voxel in $\hat{\mathcal{Y}}$ is assigned a label from the set $\mathcal{L}=\{0,1,...,C\}$, where 0 indicates empty space and $1$ to $C$ correspond to semantic classes. The model aims to infer $\hat{\mathcal{Y}}\in\mathcal{L}^{H\times W\times D}$ based on the sparse input $\mathcal{X}$, including predictions for unobserved regions, to reconstruct a complete semantic representation of the scene.
@@ III-C Center Query
$q_{c}=\begin{cases}\mathrm{FF}_{q}(f_{c}),&\text{if the}~f_{c}~\text{is non-empty}\\ \mathrm{FF}_{q}(p_{c}),&\text{otherwise}\end{cases}$ (3)
@@ V Conclusions
In this work, we propose SWA-SOP, a novel semantic occupancy prediction framework. By adopting a sliding-window mechanism that selectively computes attention over valid regions, SWA-SOP significantly improves the efficiency of global attention-based models. Within each window, we leverage spatial embeddings and queries to perform localized attention, offering greater flexibility than traditional convolutional operations. Experiments on the SemanticKITTI benchmark demonstrate that our LiDAR-based method achieves state-of-the-art performance compared to existing approaches. Furthermore, we show that the proposed SWA module can be effectively integrated into image-based pipelines, leading to consistent performance gains. These results highlight the versatility of SWA and its potential for deployment across diverse sensor configurations in autonomous driving perception systems. In future work, we plan to further integrate SWA into recent advances in cross-modal fusion [38, 39] and prompt-based representation learning [40], enabling more scalable and flexible deployment in large-scale multimodal perception systems. Additionally, to support deployment on resource-constrained platforms, we aim to incorporate lightweight modeling techniques [41, 42] to reduce computational cost without sacrificing performance.